OpenStudy (anonymous):

How do I find the instantaneous velocity at time 0.55 s? Please help!

OpenStudy (anonymous):

Time (s) 0.0 0.20 0.30 0.40 0.50 0.60 Position (cm [W]) 0 25 75 75 75 0

OpenStudy (anonymous):

I tried this: Vin = VF- Vi ------ delta time = 37.5 cm [W} - 0 -------------- 5s = 7.5 cm/s [W]

OpenStudy (anonymous):

A good approximation of the instantaneous velocity at t=55s is the average velocity from t=50s to t=60s

OpenStudy (anonymous):

So.. what happens?

OpenStudy (anonymous):

because 50s to 60s if I plot the graph is 5 s..

OpenStudy (anonymous):

The average velocity from time t=50 to t=60 is: v=(0-75)/(0.60-0.50)=-750cm/s. Because this is the average between the two times we assign it to the middle time t=55s. You can only estimate since you have discrete data

OpenStudy (anonymous):

ohh..

OpenStudy (anonymous):

'cause on the back of the book it says: 7.5 m/s [E]